OpenStudy (anonymous):

y=3x+1 , y=2x-1 solve for x and y

OpenStudy (sgadi):

x=-2 y=-5

OpenStudy (radar):

Add the two equations getting 2y=5x Solve for y, y=x5/2 Substitute this value for y x5/2=3x+1 3x-x5/2=-1 x1/2=-1 x=-2

OpenStudy (radar):

substitute -2 for x in either equation and solve for y